debra invested $18000.Part of it was invested at 3% annual simple interest, and the rest was invested at 4%. Her interest income for the first year was $650.How much did she invest at each rate?
Found 2 solutions by mananth, robertb:
Answer by mananth(16949)

debra invested $18000.Part of it was invested at 3% annual simple interest, and the rest was invested at 4%. Her interest income for the first year was $650.How much did she invest at each rate?
...
$x @ 3%
18000-x@4%
Interest = 650
...
3x+4(18000-x)=65000
3x+72000-4x=65000
-x=-7000
x= $7000 @ 3%
balance $11000 @4%
...
m.ananth@hotmail.ca

Answer by robertb(5830)

x = investment at 3%,
18,000 - x = investment at 4%. Then
0.03x + 0.04(18,000 - x) = 650.
0.03x + 720 - 0.04x = 650
-0.01x = -70
x = 7000, investment at 3%. and
18,000 - x = 11,000 = investment at 4%.